The real risk-free rate, r*, is 1.8%. Inflation is expected to average 3.5% a year for the next 4 years, after which time inflation is expected to average 4.65% a year. Assume that there is no maturity risk premium. An 8-year corporate bond has a yield of 9.15%, which includes a liquidity premium of 0.9%. What is its default risk premium? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Average Inflation for 8 years = [3.5+3.5+3.5+3.5+4.65+4.65+4.65+4.65]/8

      = 32.6 /8

       = 4.075%

Yield for 8 years =Real rate + average inflation+liquidity premium+Default risk premium

9.15 = 1.8+4.075 + .9+ DRP

9.15 = 6.775+ DRP

DRP = 9.15 -6.775

    = 2.375 OR 2.38% rounded